# EgoTextVQA: Towards Egocentric Scene-Text Aware Video Question Answering

### v1

@@ 5 Heuristic Solution Investigations
Q2: Is temporal grounding important? Sometimes. We take the three frames used in question generation as model inputs (“QA Frames”). Table 5 shows that this effectively boosts all model performances on EgoTextVQA-Indoor. However, the improvements on EgoTextVQA-Outdoor are not stable among models. Specifically, Qwen2-VL [47] and InternVL2-8B [8] increase their accuracy by 3.7% and 2.1% respectively but MiniCPM-V 2.6 [53] and Gemini 1.5 Pro decrease by 2.1% and 7.1%, respectively. Overall, the improvement suggests there is space for models to optimize their temporal grounding, i.e., to localize the key frames for answering. In contrast, the decrease in performance indicates that the models may be vulnerable to seriously reduced information input. It may also reflect the great efforts made by annotators for correcting mistakes from the automatic QA generation, as also exemplified by the poor results of GPT-4o (QA generator).
@@
Q3: Are high-resolution frames for scene text critical? Yes. Most existing MLLMs are limited to accepting fixed, low-resolution images as input. When higher resolutions are used, the number of input frames is limited. We compare the use of “Low-res. video” inputs and the specified three frames for question generation as “Low-res. frames” and “High-res. frames”. Here we choose GPT-4o [35] and three open-source models that can accept frames with resolutions of up to 1920$\times$1080 and 1280$\times$720. Figure 7 shows that answering with “QA Frames” at low resolution is comparable to uniform video sampling. Yet, with the increase of resolution, GPT-4o [35], Qwen2-VL [47], InternVL2-8B [8], and MiniCPM-V 2.6 [53] all show profound improvements, up to 18.6$\sim$31.0%.
@@ V Q ST Accuracy Score Accuracy Score
Aside from increasing the global frame resolution, we particularly study increasing the local scene-text resolution while keeping the frame resolution unchanged. To this end, we apply Microsoft Azure OCRfootnotetext: https://azure.microsoft.com/en-us/services/cognitive-services to detect scene text regions and then exploit the SOTA scene text super-resolution model DiffTSR [59] to increase the resolution of the detected scene text regions to 380 $\times$ 128. To project the super-resolution scene text back to the original low-resolution video frames, we consider three options: resize to 1.00 $\times$, 1.25$\times$, and 1.50$\times$ of the original scene text size. The results in Table 6 show that the higher resolution scene text effectively boosts the model performances on EgoTextVQA-Outdoor. Nevertheless, its improvements on EgoTextVQA-Indoor are unstable and the performances even decline with super-resolution scene text input. We speculate that scene-text detection in Ego4D videos is harder than in RoadTextVQA videos, due to lower resolution and challenging perspective. The lower performance improvements on EgoTextVQA-Indoor (vs. Outdoor) with additional OCR results in Table 7 further consolidate our speculation.
@@ 6 Conclusion
To provide a testing ground to evaluate model capacity and develop new techniques for egocentric scene-text aware QA assistance, we carefully constructed the EgoTextVQA benchmark. EgoTextVQA highlights QA assistance involving scene text from an ego-perspective in diverse real-life scenarios, including mainly outdoor driving and indoor house-keeping, with the questions reflecting the real user needs yet without the visual focus on scene text. With well-classified visual scenarios and question types, we comprehensively analyzed models that excel on existing scene-text VQA benchmarks and found that they perform poorly on EgoTextVQA. We further explore heuristic solutions and provide insights for improvements, With these efforts, we hope this work can complement existing VideoQA research in advancing real-life egocentric QA assistance.
@@ B.1 Model Details
  • InternVL2 [8] integrates InternViT-300M [9] with InternLM2.5 [5] via a randomly initialized MLP projector. It is trained on OCR datasets generated by PaddleOCR [22], utilizing Chinese images from Wukong and English images from LaionCOCO [40]. Building on the strong visual representations and high-resolution image processing capabilities of InternVL1.5 [8], InternVL2 incorporates instruction tuning, enabling competitive performance in document and chart comprehension, infographics QA, scene text understanding, OCR, and multimodal reasoning tasks.
@@ B.2 Study of MLLM Design
As the results in Tables 3 and 4, we analyze key factors contributing to the superior performance of strong models (Qwen2-VL [47], LLaVA-NeXT-Video [58], and InternVL2 [9]): (1) enhanced visual encoder capable of handling high-resolution and long-video inputs. As shown in Table 8, increasing the number of video frames and resolution improves Qwen2-VL’s performance by 1.2% and 5.8%, respectively.; (2) more powerful LLM backbones. Compared with InternVL2-8B, InternVL2-26B performance has a 7% increases; and (3) large-scale OCR training data. Beyond the commonly used TextVQA datasets, InternVL2 leverages PaddleOCR to generate additional OCR training samples. Additionally, we observe that as the number of video frames increases, Qwen2-VL’s performance improves, whereas InternVL2’s declines, underscoring the effectiveness of Qwen2-VL’s video embedding design.
@@ B.4 Human Study
The human results in Tables 3 and 4 are based on two rounds of standard human studies. Based on the reason analysis for the poor human performance in Section 4.2, we further validate the human performance by reducing the scene text recognition challenge. We sample 100 additional questions for humans to answer by providing the corresponding question frames. Table 11 shows that humans perform better without the challenge of temporal grounding but still lag behind the best closed-source model (GPT-4o [35]). This suggests advanced models may surpass humans in scene-text recognition or external knowledge, highlighting the importance of research on scene-text QA assistance.

### v2

@@ 5 Heuristic Solution Investigations
Q2: Is temporal grounding important? Sometimes. We use the three frames from question generation as model inputs (“QA Frames”), which consistently improve performance on EgoTextVQA-Indoor but have inconsistent effects on EgoTextVQA-Outdoorin Table 5. Qwen2-VL [50] and InternVL2-8B [7] increase their accuracy by 3.7% and 2.1% respectively but MiniCPM-V 2.6 [59] and Gemini 1.5 Pro decrease by 2.1% and 7.1%, respectively. Overall, the improvement suggests there is space for models to optimize their temporal grounding, i.e., to localize the key frames for answering. In contrast, the decrease in performance indicates that the models may be vulnerable to seriously reduced information input. It may also reflect the great efforts made by annotators for correcting mistakes from the automatic QA generation, as also exemplified by the poor results of GPT-4o (QA generator).
@@
Q3: Are high-resolution frames for scene text critical? Yes. Most existing MLLMs are limited to accepting fixed, low-resolution images as input. When higher resolutions are used, the number of input frames is limited. We compare the performance of the model when using the “low-resolution video” and the three key “low-resolution frames” and “high-resolution frames” specified for question generation as input. Here we choose GPT-4o [37] and three open-source models that can accept frames with resolutions of up to 1920$\times$1080 and 1280$\times$720. Figure 7 shows that answering with three key frames (“QA Frames”) at low resolution is comparable to uniform video sampling. Yet, with the increase of resolution, GPT-4o [37], Qwen2-VL [50], InternVL2-8B [7], and MiniCPM-V 2.6 [59] all show profound improvements, up to 18.6$\sim$31.0%.
Aside from increasing the global frame resolution, we focus on increasing the local scene-text resolution while keeping the global frame resolution unchanged. We use Microsoft Azure OCRfootnotetext: https://azure.microsoft.com/en-us/services/cognitive-services for scene text detection and apply the SOTA scene text super-resolution model DiffTSR [65] to increase the resolution of the detected regions to 380 $\times$ 128. To project the super-resolution scene text back to the original low-resolution video frames, we consider three options: resize to 1.00 $\times$, 1.25$\times$, and 1.50$\times$ of the original scene text size. Table 6 shows that higher-resolution scene text improves performance on EgoTextVQA-Outdoor but yields inconsistent results on EgoTextVQA-Indoor. We speculate that scene text detection in Ego4D videos is more challenging than in RoadTextVQA due to lower resolution and complex perspectives. This is further supported by the smaller performance gains on EgoTextVQA-Indoor (vs. Outdoor) with additional OCR results in Table 7.
@@ 6 Conclusion
To provide a benchmark for research on egocentric scene-text aware QA assistance, we carefully construct the EgoTextVQA dataset. EgoTextVQA highlights QA assistance involving scene text from an ego-perspective in diverse real-life scenarios, including outdoor driving and indoor house-keeping. The questions reflect the real user needs, yet the visual attention is often not focused on the scene text. With well-classified visual scenarios and question types, we comprehensively analyze models that excel on existing scene-text VQA benchmarks and find that they struggle on EgoTextVQA. We further explore heuristic solutions and provide insights for improvements, With these efforts, we hope this work can complement existing VideoQA research in advancing real-life egocentric QA assistance.
7 Acknowledgements
This research was funded by the NUS Artificial Intelligence Institute (NAII) seed grant number NAII-SF-2024-003. Additionally, this work was supported by the National Key R&D Program of China (NO.2024YFB3311602), the National Natural Science Foundation of China (72188101, 62020106007, 62272435), the Major Project of Anhui Province (202203a05020011), the Anhui Provincial Natural Science Foundation (2408085J040), the Fundamental Research Funds for the Central Universities (JZ2023YQTD0072, JZ2024HGTG0309, and JZ2024AHST0337), and the New Cornerstone Science Foundation through the XPLORER PRIZE. We sincerely appreciate the dedication and meticulous efforts of our dataset annotators, whose contributions were invaluable in ensuring data quality and advancing this research.
@@ B.1 Model Details
  • InternVL2-8B [7] integrates InternViT-300M [8] with InternLM2.5-7B [5] via a randomly initialized MLP projector. It is trained on OCR datasets generated by PaddleOCR [22], utilizing Chinese images from Wukong and English images from LaionCOCO [42]. Building on the strong visual representations and high-resolution image processing capabilities of InternVL1.5 [7], InternVL2 incorporates instruction tuning, enabling competitive performance in document and chart comprehension, infographics QA, scene text understanding, OCR, and multimodal reasoning tasks.
@@ B.2 Study of MLLM Design
We analyze key factors contributing to the superior performance of strong models (Qwen2-VL [50], LLaVA-NeXT-Video [64], and InternVL2 [8]): (1) enhanced visual encoder capable of handling high-resolution and long-video inputs. In Table 8, increasing the number of video frames and resolution improves Qwen2-VL’s performance by 1.2% and 5.8%; (2) more powerful LLM backbones. Compared with InternVL2-8B, InternVL2-26B performance has a 7% increases in Table 8; and (3) large-scale OCR training data. Beyond the commonly used TextVQA datasets, InternVL2 leverages PaddleOCR to generate OCR samples for training. Additionally, we observe that as the number of video frames increases, Qwen2-VL’s performance improves, whereas InternVL2’s declines, underscoring the effectiveness of Qwen2-VL’s video embedding design.
@@ B.4 Human Study
In Section 4, the human results are based on two rounds of standard human studies. Based on the reason analysis for the poor human performance in Section 4.2, we further validate the human performance by reducing the scene text recognition challenge. We sample 100 additional questions for humans to answer by providing the corresponding question frames. Table 11 shows that humans perform better without the challenge of temporal grounding but still lag behind the best closed-source model (GPT-4o [37]). This suggests advanced models may surpass humans in scene-text recognition or external knowledge, highlighting the importance of research on scene-text QA assistance.